An investor purchased 534 shares of common stock $24 par for $26,166 subsequently 92 shares were sold for $27 per share what is the amount of gain or loss on the sale


1
Expert's answer
2021-10-25T17:41:33-0400

Sales proceeds of shares"=92\times27=\$2484"

less: cost of shares"=92shares \times 26166/534per \space share=92\times 49=\$4508"

Gain on sale of shares"=4508-2484=\$2024"